# On Fairness of Task Arithmetic: The Role of Task Vectors

### v1

@@ Abstract
Model editing techniques, particularly task arithmetic using task vectors, have shown promise in efficiently modifying pre-trained models through arithmetic operations like task addition and negation. Despite computational advantages, these methods may inadvertently affect model fairness, creating risks in sensitive applications like hate speech detection. However, the fairness implications of task arithmetic remain largely unexplored, presenting a critical gap in the existing literature. We systematically examine how manipulating task vectors affects fairness metrics, including Demographic Parity and Equalized Odds. To rigorously assess these effects, we benchmark task arithmetic against full fine-tuning, a costly but widely used baseline, and Low-Rank Adaptation (LoRA), a prevalent parameter-efficient fine-tuning method. Additionally, we explore merging task vectors from models fine-tuned on demographic subgroups vulnerable to hate speech, investigating whether fairness outcomes can be controlled by adjusting task vector coefficients, potentially enabling tailored model behavior. Our results offer novel insights into the fairness implications of model editing and establish a foundation for fairness-aware and responsible model editing practices.
On Fairness of Task Arithmetic: The Role of Task Vectors
Hiroki Naganumathanks: Equal contributions. thanks: Correspondence: naganuma.hiroki@mila.quebec Université de Montréal, Mila                        Kotaro Yoshida11footnotemark: 1 Institute of Science Tokyo                        Laura Gomezjurado Gonzalez11footnotemark: 1 Stanford University
Takafumi Horie Kyoto University                        Yuji Naraki Independent Researcher                        Ryotaro Shimizu University of California San Diego, ZOZO Research
@@ 1 Introduction
As large language models (LLMs) see broader application, efficient techniques for adapting them to specific tasks become increasingly crucial. While some models have been distilled (Sanh et al., 2019; Jiao et al., 2020; Turc et al., 2020) or are relatively small (Abdin et al., 2024), task-specific fine-tuning often demands substantial computational resources, prompting the development of parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2022; Ben Zaken et al., 2022; Dettmers et al., 2023).
@@
One notable example is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which updates a compact set of parameters while leaving most of the original weights untouched, thus reducing training costs. Despite the popularity of PEFT methods, they do not resolve every challenge: in high-stakes tasks with imbalanced data, LoRA and similar approaches can preserve or even amplify biases, raising concerns about fairness (Ding et al., 2024b; Sap et al., 2019).
An alternative strategy that has recently drawn attention is model editing with task vectors (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025). A task vector is defined as the parameter difference between a base pre-trained model $\theta_{\mathrm{base}}$ and a fine-tuned model $\theta_{\mathrm{task}}$. By adding or subtracting this vector within the original weight space (so-called “task arithmetic”), a user can edit or remove the corresponding task-specific behavior without further gradient-based training (Ilharco et al., 2023). Moreover, scaling the task vector grants fine-grained control over the strength of the transferred capability. This approach represents a promising direction, as it directly manipulates parameters while avoiding a costly re-optimization of the entire model.
In addition to these computational benefits, prior work has suggested that separating and analyzing task vectors may enhance interpretability (Cerrato et al., 2025). By isolating the weight updates associated with particular subgroups (e.g., racial or gender demographics), one can potentially trace how the model adapts to each subgroup. This feature is appealing for investigating biases arising from unequal representation in training data, as it highlights which groups require larger shifts in weight space. Nevertheless, open questions persist regarding how well this model editing using task-vector preserves or exacerbates fairness. For instance, improving performance for one demographic might degrade outcomes for another, and it is not yet clear how to balance trade-offs with established fairness metrics such as Demographic Parity (DPD) or Equalized Odds (EOD).
To address this gap, we systematically examine how task-vector editing compares to both traditional full-parameter fine-tuning (FFT) and LoRA, and we further explore whether injecting task vectors into an FFT model offers additional control over fairness. Our experiments focus on hate-speech detection on Llama-7B (Touvron et al., 2023) , measured by subgroup-specific accuracy and widely used fairness metrics. Our contributions and findings are summarized as follows:
@@
  • A thorough comparison of four algorithms (FFT, LoRA, model editing using task-vector, and a hybrid approach injecting task vectors into FFT) in terms of their effects on fairness metrics and overall performance (Figure  2)].
  • An analysis showing that task vectors can substantially improve fairness while preserving accuracy, provided that their scalar coefficients are appropriately tuned (Figure 3).
  • Evidence that merging task vectors for underrepresented subgroups with existing models can adjust fairness outcomes without incurring a significant accuracy drop (Figures 4(a), 4(b) and  5(a)).
Through this analysis, we illustrate how task vectors can reduce risks from a fairness perspective while taking advantage of their flexibility and interpretability as a model editing approach. These findings provide a foundation for extending task-vector-based methods to promote fair and responsible operation of large language models.
@@
Figure 2: LoRA and FFT vs. Task addition with the optimal coefficient for the training accuracy ($\lambda=0.8$ for gender setting and $\lambda=0.5$ for race setting) on group-wise accuracy, demographic parity difference (DPD, lower is fairer), and equalized odds difference (EOD, lower is fairer). Error bars denote the standard error across three seeds. Columns: group-wise accuracy, DPD, EOD. No consistent pattern emerges that task addition necessarily degrades subgroup fairness relative to LoRA or FFT subgroups show improvements or comparable results under task addition, while others show small declines.
@@ 2 Preliminaries
Negation: Using the negative of a task vector, $-\Delta\theta$, one can subtract a task’s influence. For example, applying $\theta_{\text{base}}-\Delta\theta_{A}$ (or equivalently $\theta_{\text{base}}+(-\Delta\theta_{A})$) yields a model with reduced performance on task A—effectively unlearning or forgetting it—while preserving other behaviors (Ilharco et al., 2023). This is useful for removing undesirable skills or biases.
Scalar scaling: Multiplying a task vector by a scalar $\lambda$ adjusts the strength of the edit. For example, using $\theta_{\text{base}}+\lambda\Delta\theta_{A}$ allows partial ($10<\lambda<$ 1) or amplified ($\lambda>1$) application of a task’s effect. This scaling provides fine-grained control over how strongly the task knowledge is injected into the model.
@@
Since task vectors reside in a common weight space, they can be merged by simple addition with tunable scaling. Formally, given a base model $\theta_{0}$ and task vectors $\Delta\theta_{i}$, one can construct a merged model as:
$\theta_{\text{merged}}=\theta_{0}+\sum_{i}\lambda_{i}\,\Delta\theta_{i}\,,$ (1)
where each coefficient $\lambda_{i}$ controls the influence of task $i$. Varying $\lambda_{i}$ thus directly modulates how strongly the $i$-th task’s knowledge is injected, allowing fine-grained blending of capabilities. Indeed, adding multiple task vectors with $\lambda_{i}=1$ endows a model with all those capabilities simultaneously (Ilharco et al., 2023). Optimizing the $\lambda_{i}$ values (i.e., learning an anisotropic scaling for each vector) further improves the composition by balancing contributions and reducing interference between tasks (Zhang et al., 2024).
@@ 3 Related Work
Task arithmetic: efficiency and interpretability.
Task vectors offer a computationally efficient framework for editing and analyzing model behavior. Once a task vector is computed—namely, the weight difference between a base model and its fine-tuned variant (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025)—no additional training data or retraining is required to transfer or remove task-specific capabilities. By treating each fine-tuning update as a direction in weight space, practitioners can combine or negate these updates through simple addition or subtraction (Ilharco et al., 2023). This modularity not only reduces computational overhead but also enhances interpretability by isolating the contribution of each task.
Beyond modularity, task arithmetic can reveal valuable information about how and where a model adapts to new tasks. Li et al. (2024) show a near-linear relationship between data size and the norm of a task vector, suggesting that over-represented tasks can dominate weight space shifts in multi-task settings. In addition, the orientation of task vectors can indicate synergies or conflicts among tasks (Li et al., 2025), and decomposing these vectors by layer can pinpoint which parts of the model are most affected (Zhang et al., 2024; Gargiulo et al., 2025). Hence, task vectors offer a promising lens for diagnosing training dynamics and identifying potential biases.
Fairness metrics for LLMs.
Fairness in large language models is commonly evaluated using criteria such as Demographic Parity, Equalized Odds, and accuracy parity. Demographic Parity requires similar positive outcome rates across demographic groups, while Equalized Odds demands that true and false positive rates be equivalent. Accuracy parity checks for consistent predictive performance across groups (Fraenkel, 2020; Kennedy et al., 2020a; Pitoura, 2019; Quan et al., 2023). These metrics are broadly used to detect biases and measure whether a model’s behavior disproportionately disadvantages certain populations.
@@
Parameter-efficient methods such as LoRA (Hu et al., 2022) address computational bottlenecks by training only a small set of parameters, yet they do not inherently solve fairness issues. In some cases, LoRA yields comparable subgroup performance to full fine-tuning (Ding et al., 2024b), while in others, it fails to mitigate toxic behaviors or biases (Das et al., 2024). The variance in outcomes depends on factors like the rank of the LoRA matrices, the base model’s quality, and the distribution of training data  (Das et al., 2024).
Merging tasks and fairness considerations.
@@
Despite the potential efficiency gains and interpretability offered by task arithmetic, the merging of task vectors for multiple groups can trigger new challenges. For instance, simply summing vectors may lead to “negative transfer,” where updates beneficial to one subgroup degrade performance for another (Ding et al., 2024a; Yu et al., 2020). In highly imbalanced settings, merging models through supervised fine-tuning can also disproportionately favor majority groups while disadvantaging minorities (Cross et al., 2024). Because fairness does not compose additively, interactions among subgroup-specific task vectors can produce unpredictable shifts in metrics like Demographic Parity and Equalized Odds (Gohar et al., 2023).
@@
Consequently, identifying effective ways to adjust task vectors—such as through scalar scaling—remains a key step toward fairness-aware model editing. This work aims to fill that gap by systematically evaluating how these operations influence both fairness and overall model accuracy.
4 Experimental Setup
4.1 Configuration.
Building on the experimental framework established by Ding et al. (2024b), we adopted their evaluation and experimental procedure to assess the fairness implications of LoRA in comparison to FFT. In our work, we extend this analysis by focusing on how task arithmetic compares to both LoRA and FFT in terms of fairness and performance. The detailed experimental setup is provided in Appendix B.
@@ 4 Experimental Setup
Table 1: Data statistics in the gender and race subgroups.
Dataset.
We use a modified version of the Berkeley D-Lab Hate Speech dataset originally introduced by Kennedy et al. (2020a) and adapted by Ding et al. (2024b), the research we are building upon. Our dataset contains a total of 6,898 tweet-sized text snippets annotated for hate speech and categorized by sensitive attributes: Race and Gender, each further divided into fine-grained subgroups (e.g., Women, Non-binary, Men within Gender) as shown in Table 1. We frame hate speech detection as a binary classification task: given a text snippet, the model predicts whether it constitutes hate speech (e.g., hatespeech in the Gender subset may target Non-binary or Trans Women). Each example includes both the hate speech label and one or more protected attribute annotations (e.g., gender = woman, race = Asian). These are used to assess subgroup-level performance and fairness metrics.
This setting supports rigorous fairness analysis due to its rich attribute annotations and real-world relevance. Hate speech detection is a challenging and high-stakes classification problem: it requires models to identify subtle or implicit harm, resolve linguistic ambiguity, and perform robustly across diverse dialects and identity references (Kennedy et al., 2020a). As models increasingly mediate content moderation, ensuring reliable and equitable hate speech detection is essential for safe deployment in real-world systems.
Evaluation metrics.
We evaluate each method on both predictive performance and fairness metrics. Our goal is to understand how scaling or merging task vectors affects these measures.
  • Predictive Performance:
    Accuracy:
    Standard metric for classification tasks, measuring the percentage of correct predictions.
  • Fairness Metrics: We also adopt the metrics used for Ding et al. (2024b) to quantify disparate performance across protected subgroups. These metrics are widely used for fairness research in ML:
    Demographic Parity Difference (DPD):
    Measures the disparity in the model’s positive prediction rates across sensitive attribute groups. A smaller DPD indicates that the model assigns positive outcomes at similar rates across these groups, reflecting a more uniform treatment irrespective of group membership (Agarwal et al., 2018, 2019).
    Equalized Odds Difference (EOD)
    : Measures the disparity in the model’s true and false positive rates across sensitive attribute groups. A smaller EOD indicates that the model’s overall error rate is more balanced across groups (Das et al., 2024).
4.2 Protocol.
We evaluate our methods using a main base model: LLaMA2-7B [^1] LLaMA 2 is licensed under the LLAMA 2 Community License, Copyright (c) Meta Platforms, Inc. All Rights Reserved. See: https://ai.meta.com/llama/license. Our fairness evaluations focus on two sensitive attributes: gender and race, using subgroup-wise metrics mentioned earlier –accuracy, DPD, and EOD.

### v2

@@ Abstract
Model editing techniques, particularly task arithmetic with task vectors, offer an efficient alternative to full fine-tuning by enabling direct parameter updates through simple arithmetic operations. While this approach promises substantial computational savings, its impact on fairness has remained largely unexplored—despite growing concern over biased outcomes in high-stakes applications such as hate speech detection. In this work, we present the first systematic study of group fairness in task arithmetic within this binary text and image classification regime, comparing it against full fine-tuning (FFT) and Low-Rank Adaptation (LoRA). We evaluate across multiple language models and datasets using standard group fairness metrics, including Demographic Parity and Equalized Odds. Our analysis shows that task vectors can be tuned to achieve competitive accuracy while reducing disparities, and that merging subgroup-specific task vectors provides a practical mechanism for steering fairness outcomes. We further provide a theoretical bound linking task vector scaling to fairness metrics, offering insight into the observed trade-offs. Together, these findings establish task arithmetic not only as a cost-efficient editing method but also as a fairness-aware alternative to existing adaptation techniques, within the standard group-fair classification setting, laying the groundwork for responsible deployment of large language models. Our code is available at https://github.com/LauraGomezjurado/fairness_task_vector_deploy
[…]
@@ 1 Introduction
As large language models (LLMs) are deployed across increasingly diverse applications, efficient techniques for adapting them to specific tasks have become essential. While model distillation and compact architectures reduce computational demands (Sanh et al., 2019b; Jiao et al., 2020; Turc et al., 2020; Abdin et al., 2024), task-specific fine-tuning (FFT) remains resource-intensive. This has motivated parameter-efficient fine-tuning (PEFT) methods such as adapters and Low-Rank Adaptation (LoRA) (Houlsby et al., 2019; Hu et al., 2022; Ben Zaken et al., 2022; Dettmers et al., 2023), which update only a small fraction of parameters.
@@
LoRA exemplifies this trade-off: it preserves most of the pretrained weights while reducing training costs. However, PEFT methods do not resolve deeper concerns. In high-stakes domains with imbalanced data—such as toxicity or hate-speech detection—they can maintain or even amplify biases (Ding et al., 2024; Sap et al., 2019), raising concerns about fairness.
A promising alternative is task arithmetic with task vectors (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025b; Yoshikawa et al., 2025; Yoshida et al., 2025a). A task vector is defined as the difference between a fine-tuned model and its base counterpart. By adding, subtracting, or scaling such vectors, one can directly edit model behavior without gradient-based retraining. This approach offers (i) computational efficiency, (ii) fine-grained control over transferred capabilities, and (iii) enhanced interpretability when task vectors are associated with specific subgroups (Cerrato et al., 2022). Yet its fairness implications remain poorly understood. For example, enhancing performance on one demographic subgroup may inadvertently degrade outcomes for another, and the trade-offs with standard metrics such as Demographic Parity (DPD) or Equalized Odds (EOD) remain unclear.
To address this gap, we conduct the first systematic study of fairness in task arithmetic. We compare task vector editing against both FFT and LoRA, and we further investigate whether injecting subgroup-specific task vectors into an FFT model provides additional control over fairness outcomes. In the NLP domain, our experiments focus on hate-speech detection with LLaMA-7B (Touvron et al., 2023) and toxicity detection on Civil Comments (Borkan et al., 2019) using DistilBERT and Qwen2.5-0.5B (Qwen Team, 2025). In the computer-vision domain, we evaluate age classification on UTKFace (Zhang et al., 2017) using ViT-Base/16 (Dosovitskiy et al., 2021). Across both domains and model architectures, we observe consistent fairness–utility trade-offs.
Our contributions are as follows:
@@
  • Comprehensive evaluation: We compare FFT, LoRA, task vector editing, and a hybrid approach that injects task vectors into FFT, analyzing their impact on fairness metrics and predictive performance (Figure 2).
  • Fairness through scaling: We show that adjusting task vector coefficients can substantially improve fairness while maintaining accuracy (Figure 3).
  • Subgroup-sensitive editing: We demonstrate that merging task vectors from underrepresented subgroups allows targeted fairness adjustments with negligible accuracy loss (Figures 4(a), 4(b), 5(a)).
  • Theoretical grounding: We derive an upper bound linking task vector scaling to DPD and EOD, providing a principled explanation for the observed fairness–accuracy trade-offs (Section Theorem (informal) and Appendix C).
Throughout, the prediction task remains binary, but the fairness structure is not: we evaluate disparities across multiple demographic groups, following widely used group-fairness benchmarks (Hardt et al., 2016; Agarwal et al., 2018; Zafar et al., 2017; Kearns et al., 2018; Das et al., 2024; Dutt et al., 2023; Sukumaran et al., 2024). Taken together, our results provide an empirical foundation for understanding fairness–accuracy trade-offs in task arithmetic within the standard group-fair binary-classification regime, while pointing toward extensions to more complex tasks (e.g., multi-label or generative settings) as future work.
@@
Figure 2: LoRA and FFT vs. Task addition with the optimal coefficient for the training accuracy ($\lambda=0.8$ for gender setting and $\lambda=0.5$ for race setting) on group-wise accuracy, demographic parity difference (DPD, lower is fairer), and equalized odds difference (EOD, lower is fairer). Error bars denote the standard error across three seeds. Columns: group-wise accuracy, DPD, EOD. No consistent pattern emerges indicating that task addition systematically degrades subgroup fairness relative to LoRA or FFT. While some subgroups show improvements or comparable results under task addition, others exhibit small declines.
@@ 2 Preliminaries
Negation: Using the negative of a task vector, $\Delta\theta$, one can subtract a task’s influence. For example, applying $\theta_{\text{base}}-\Delta\theta_{A}$ (or equivalently $\theta_{\text{base}}+(-\Delta\theta_{A})$) yields a model with reduced performance on task A—effectively unlearning or forgetting it—while preserving other behaviors (Ilharco et al., 2023). This is useful for removing undesirable skills or biases.
Scalar scaling: Multiplying a task vector by a scalar $\lambda$ adjusts the strength of the edit. For example, using $\theta_{\text{base}}+\lambda\Delta\theta_{A}$ allows partial ($0<\lambda<1$) or amplified ($\lambda>1$) application of a task’s effect. This scaling provides fine-grained control over how strongly the task knowledge is injected into the model.
@@
Because task vectors live in a common weight space, they can be merged by linear combination. Let $\theta_{0}$ denote a base model and let $\{\Delta\theta_{i}\}_{i=1}^{K}$ be task vectors. A generic merged model takes the form
$\theta_{\text{merged}}=\theta_{0}+\sum_{i=1}^{K}\lambda_{i}\Delta\theta_{i},$ (1)
where the coefficients $\lambda_{i}$ govern the relative contribution of each task (Zhang et al., 2024). Prior work explores learning these coefficients with, for example, Bayesian optimization or multi-objective search over $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{K})$, but such methods introduce additional computational cost and stochasticity.
In this work we deliberately adopt the simpler and widely used single-scalar parameterization, in which a global coefficient $\lambda$ scales the entire task vector update. Concretely, the models we evaluate can be written as
$\theta(\lambda)=\theta_{0}+\lambda\,\Delta\theta,$ (2)
where $\Delta\theta$ is a fixed task vector (or sum of task vectors), and the same scalar $\lambda$ is applied uniformly across all inputs, demographic subgroups, and fairness metrics. This single global coefficient mirrors how task-arithmetic toolkits are exposed in practice, and it provides a one-dimensional, easily interpretable control knob for tracing fairness–accuracy trade-offs. We treat $\lambda$ as a hyperparameter and select it by grid search over a fixed set of candidate values on a held-out validation set, using a common objective that balances accuracy and group-fairness measures. The same validation protocol is used to tune the corresponding regularization weights for our SFT, LoRA, and task-arithmetic models, ensuring that all methods are compared under matched fairness–accuracy trade-offs.
@@ 3 Related Work
This section outlines how FFT, LoRA, task merging, and fairness are related. Further discussion of task-arithmetic efficiency, interpretability, and fairness in VLMs appears in Appendix A.
@@
Parameter-efficient methods such as LoRA (Hu et al., 2022) address computational bottlenecks by training only a small subset of parameters, yet they do not inherently resolve fairness concerns. Empirical studies show that LoRA’s impact on subgroup performance is mixed: in some settings it achieves results comparable to full fine-tuning (Ding et al., 2024), while in others it preserves undesirable biases or fails to mitigate toxic behaviors (Das et al., 2024). These divergent outcomes depend on factors such as the rank of the LoRA updates, the base model’s pre-existing biases, and the distributional properties of the fine-tuning data (Das et al., 2024). Recent work begins to incorporate fairness objectives directly into PEFT. Wang and Demberg (Wang and Demberg, 2024) introduce a multi-objective method that jointly targets accuracy and reduced stereotypical bias, and Sukumaran et al. (Sukumaran et al., 2024) propose FairLoRA, which embeds fairness-driven structure into low-rank updates for vision models. These efforts show that fairness-aware PEFT is feasible but often requires custom objectives and careful tuning.
Merging tasks and fairness composition.
@@
Despite the potential efficiency gains and interpretability offered by task arithmetic, the merging of task vectors for multiple groups can trigger new challenges. For instance, simply summing vectors may lead to “negative transfer,” where updates beneficial to one subgroup degrade performance for another (Ding et al., 2023; Yu et al., 2020). In highly imbalanced settings, merging models through supervised fine-tuning can also disproportionately favor majority groups while disadvantaging minorities (Cross et al., 2024).
@@
Additionally, prior work shows that fairness guarantees often do not compose: even if individual components satisfy group or individual fairness in isolation, composing them can break those guarantees (Dwork and Ilvento, 2018). This motivates our focus on post‑hoc task‑arithmetic edits: adding or scaling subgroup task vectors can be viewed as composing behaviors, and interactions among subgroup-specific task vectors can produce unpredictable shifts in metrics like Demographic Parity and Equalized Odds (Gohar et al., 2023). Consequently, identifying effective ways to adjust task vectors—such as through scalar scaling—remains a key step toward fairness-aware model editing. This work aims to fill that gap by systematically evaluating how these operations influence both fairness and overall model accuracy.
In parallel, multi‑task fairness methods such as Multi‑Task‑Aware Fairness (Wang et al., 2021), Learning‑to‑Teach Fairness‑Aware MTL (Roy and Ntoutsi, 2022), and FairBranch (Roy et al., 2024) manage fairness–accuracy trade‑offs during training. Our study complements these by asking: when we edit models after training via task vectors, can simple controls (e.g., $\lambda$‑scaling) recover fairer behavior without retraining?
Group fairness metrics in binary text classification.
Research on group fairness in text classification predominantly adopts the binary prediction setting (e.g., toxic vs. non-toxic) with multiple demographic groups (Hardt et al., 2016; Agarwal et al., 2018; Zafar et al., 2017; Kearns et al., 2018; Das et al., 2024; Dutt et al., 2023; Sukumaran et al., 2024). Within this framework, two metrics have emerged as canonical due to their interpretability and broad adoption: Demographic Parity Difference (DPD), which measures disparities in positive prediction rates, and Equalized Odds Difference (EOD), which measures disparities in true- and false-positive rates (Hardt et al., 2016; Feldman et al., 2015; Kennedy et al., 2020a). These metrics form the standard baseline in fairness auditing toolkits (Bellamy et al., 2018; Fairlearn contributors, 2025) and underpin much of the fair-PEFT literature (Fraenkel, 2020; Pitoura, 2019; Quan et al., 2023; Ding et al., 2024). Their centrality is motivated by theory: for calibrated, score-based classifiers with group-agnostic thresholds, many group-fairness desiderata collapse to constraints on selection-rate and error-rate disparities, and classical impossibility results show that calibration, selection parity, and error parity cannot be simultaneously satisfied when base rates differ (Kleinberg et al., 2017). Consequently, DPD and EOD reveal the essential structure of fairness–accuracy trade-offs without invoking additional structural or counterfactual assumptions.
Our evaluation follows this binary prediction framework but focuses on a multi-group setting. We report both macro-averaged and worst-group versions of DPD and EOD. This choice allows us to capture disparities across the full distribution of subpopulations, where fairness–accuracy tensions are often most pronounced. Fairness metrics for multiclass, multi-label, or generative models remain active research areas, and no widely accepted standard analogous to DPD/EOD currently exists. Given this landscape, the binary multi-group regime offers a well-defined and empirically robust foundation for systematic fairness analysis. Formal metric definitions appear in Appendix B, and implementation details are provided in §4.1.
@@
Table 1: Berkeley D-Lab Hate Speech data statistics in the gender and race subgroups.
4 Experimental Setup
4.1 Configuration
Building on the experimental framework established by Ding et al. (2024), we adopted their evaluation and experimental procedure to assess the fairness implications of LoRA in comparison to FFT. In our work, we extend this analysis by focusing on how task arithmetic compares to both LoRA and FFT in terms of fairness and performance. The detailed experimental setup is provided in Appendix D.
Datasets.
We build on the Berkeley D-Lab Hate Speech dataset introduced by Kennedy et al. (2020a) and adapted by Ding et al. (2024). Our version contains 6,898 tweet-length snippets annotated for hate speech and two sensitive attributes, Race and Gender, each with fine-grained subgroups (e.g., Women, Non-binary, Men for Gender; see Table 1). We treat hate speech detection as binary classification and use the subgroup annotations (e.g., gender = woman, race = Asian) to compute subgroup-level performance and fairness metrics. To test generalization beyond hate speech, we also evaluate on the Civil Comments dataset (Borkan et al., 2019), a large-scale toxicity corpus with sensitive-attribute labels. We binarize toxicity using a threshold of $0.5$, treating comments above this value as positive (“flagged”), and assess fairness across Gender and Race subgroups. For the vision domain, we follow Ding et al. (2024) and use the UTKFace face image dataset (Zhang et al., 2017), which provides age (in years), gender, and coarse-grained race labels. We convert it into a binary age-classification task by thresholding age at 30 years into “younger” ($\leq$30) vs. “older” (>>30), with the older group as the positive class. We treat race and gender as sensitive attributes and evaluate subgroup performance across the standard UTKFace race categories (White, Black, Asian, Indian, Others) and gender categories (Male, Female), mirroring the subgroup-based analysis in Ding et al. (2024).
Evaluation metrics and fairness scope.
Our fairness analysis follows the standard group-fairness framework for supervised classification, where two criteria form the canonical basis for auditing disparate impact across demographic groups. DPD captures disparities in selection rates, and EOD captures disparities in class-conditional error rates (TPR/FPR). These metrics are widely used, theoretically grounded, and highly interpretable: together they summarize the allocation and error-rate harms that motivate much of the fairness literature. This makes the binary text and image classification setting—where selection and error rates are well-defined—the natural domain in which to study fairness effects of task vector editing. For each protected attribute (e.g., gender, race/ethnicity), we compute subgroup-resolved DPD, EOD, and accuracy. We report per-subgroup values, macro-averages, and worst-group results. These choices mirror established practice and enable direct comparison to prior PEFT–fairness evaluations discussed in §3. Formal definitions and computation details appear in Appendix B.
4.2 Protocol
We evaluate our methods on a main generative base model, LLaMA2-7B (Touvron et al., 2023) [^1] LLaMA 2 is licensed under the LLAMA 2 Community License, Copyright (c) Meta Platforms, Inc. All Rights Reserved. See https://ai.meta.com/llama/license., and two compact baselines for CivilComments toxicity, DistilBERT (Sanh et al., 2019a) [^2] DistilBERT is released under the Apache 2.0 License. See https://github.com/huggingface/transformers/blob/main/LICENSE. and Qwen2.5-0.5B (Qwen Team, 2025) [^3] See the Qwen2.5-0.5B model card and license details at https://huggingface.co/Qwen/Qwen2.5-0.5B.. For the vision experiments on UTKFace, we use ViT-Base/16 (Dosovitskiy et al., 2021) [^4] Original ViT implementation is released under the Apache 2.0 License. See https://github.com/google-research/vision_transformer., a 12-layer Vision Transformer with 768-dimensional hidden size and 16×16 patch embeddings. We select ViT-Base/16 as it represents a standard, high-capacity encoder architecture commonly used in fairness studies (Ding et al., 2024) and provides a strong counterpart to our text-based models, enabling cross-domain comparisons. Our fairness evaluations focus on two sensitive attributes: gender and race across both datasets, computing subgroup accuracy, DPD, and EOD. These selections span distinct architectures (decoder-only vs. encoder-only), parameter scales, tasks, and label taxonomies.